The intensity distribution of solar radiation peaks at a wavelength of approximately 500 nm. Assuming the Sun radiates like a black body, calculate the flux of photons leaving the sun in W/m2.
1
Expert's answer
2020-07-22T06:03:17-0400

"R_e=\sigma T^4=5.67\cdot 10^{-8}\cdot (\frac{b}{\lambda})^4=5.67\cdot 10^{-8}\cdot (\frac{0.002898}{500\cdot10^{-9}})^4="


"\approx64\cdot10^6" "W\/m^2"